3. (20pts.) Find the Fourier series of the function given 0- <x<0 x. 0<x<
Answer #1

Sal? & fcm) = 2 0, - TT < X<o 1 x, 0<x< TT series of the function f(x) is given by - Fourier n=1 -1 An = Cos nxdx The f(x) =an = .(ท. i sin nx here x sin (na On nis an int an= *[fo.comadx + 5*x con ca] an= 2 px x cosne dox { Apply Integration by par, Now, on = 0 {cos na = 60} f() = + ( cost and coone + Cung on in nx) fony = 4 + leyti. Co nx + cyn din na) Ans.